OpenStudy (anonymous):

What is the new volume of a rectangular prism that is originally one hundred ninety-two cubic meters and is then resized using a scale factor of one-quarter on each side length? three cubic meters eight cubic meters twelve cubic meters forty-eight cubic meters

OpenStudy (anonymous):

i think it is 12

OpenStudy (campbell_st):

well if you look at it this way original has 4 to 1 in linear... so volume ratios will be 4^3 : 1^3 so 64 to 1 the you can either use ratios 1/64 = V/192 and solve for V the volume of the smaller or just find 1/64 of the original volume... basically the same method hope it helps

OpenStudy (anonymous):

so do i just cross multiply?

OpenStudy (campbell_st):

yep either way its V = 1/64*192

OpenStudy (anonymous):

thanks :)